### v1

# Neural Scaling Laws on Graphs

Abstract
Deep graph models (e.g., graph neural networks and graph transformers) have become important techniques for leveraging knowledge across various types of graphs. Yet, the scaling properties of deep graph models have not been systematically investigated, casting doubt on the feasibility of achieving large graph models through enlarging the model and dataset sizes. In this work, we delve into neural scaling laws on graphs from both model and data perspectives. We first verify the validity of such laws on graphs, establishing formulations to describe the scaling behaviors. For model scaling, we investigate the phenomenon of scaling law collapse and identify overfitting as the potential reason. Moreover, we reveal that the model depth of deep graph models can impact the model scaling behaviors, which differ from observations in other domains such as CV and NLP. For data scaling, we suggest that the number of graphs can not effectively metric the graph data volume in scaling law since the sizes of different graphs are highly irregular. Instead, we reform the data scaling law with the number of edges as the metric to address the irregular graph sizes. We further demonstrate the reformed law offers a unified view of the data scaling behaviors for various fundamental graph tasks including node classification, link prediction, and graph classification. This work provides valuable insights into neural scaling laws on graphs, which can serve as an essential step toward large graph models.
Machine Learning, ICML
1 Introduction
Neural scaling laws (Rosenfeld et al., 2019; Kaplan et al., 2020; Abnar et al., 2021) quantitatively describe how model performance grows with the scale of training. There are two basic forms: data scaling law (how the performance changes with training dataset size) and model scaling law (how the performance changes with model size). Neural scaling laws guide the construction of large models from two perspectives: (1) They enable us to predict model performance on large-scale training by extrapolating from small-scale runs with lower computational budgets. For instance, the performance of GPT-4 on coding problems fits accurately with the prediction of the scaling law (Achiam et al., 2023). (2) They enable us to identify better model architectures through small-scale training, typically selecting those with the largest performance gains for a given computational overhead. Such laws have played an instrumental role in the recent successes of large models in natural language processing (NLP) and computer vision (CV) domains (Brown et al., 2020; Touvron et al., 2023; Chowdhery et al., 2023; Zhai et al., 2022; Dehghani et al., 2023).
Despite these impacts on NLP and CV, no similar revolution has occurred in the graph machine learning domain (Zhang et al., 2023), given its unique difficulties. The investigation of neural scaling laws on graphs faces new challenges from both data and model perspectives. (1) From the data perspective: Graphs are irregular and diverse. Both the number of nodes and edges of a graph can vary dramatically. As an example, in the ogbg-ppa dataset of $158k$ graphs (Szklarczyk et al., 2019), the number of nodes in each graph ranges from $50$ to $300$, and the number of edges from $54$ to $18,096$, indicating significant variability across the dataset. The variability in size suggests that the current metric (graphs) insufficiently represents the data volume, hinting at the need for a new graph scaling law metric. (2) From the model perspective: Besides the model weights considered in the existing scaling law, deep graph models also need to consider the influence of the non-parametric aggregations, i.e., whether graph models with different numbers of aggregation layers could exhibit distinct scaling behaviors. This issue is not yet examined.
In this work, we make a first attempt to address these unique challenges and investigate the model scaling law and the data scaling law on graphs. While the main focus of this work is on the graph classification tasks, we also show that our settings and results can be extrapolated to node classification and link prediction tasks. Experimental observations reveal neural scaling laws across different models and dataset sizes for graph learning tasks. Our contributions can be summarized as follows:
  • For the model scaling law: (i) We examine the model scaling law for graph learning, which reveals that performance scales with model size, but collapses when the model becomes excessively large, likely due to overfitting. This suggests graph learning models are likely to be data hungry. (ii) From the perspective of architecture, we find that deep graph models of different numbers of layers will exhibit different model scaling behaviors, contrasting with the conclusions in CV and NLP domains.
  • For the data scaling law: (i) Regarding the size variability problem, we reveal that the number of edges is a better metric than the number of graphs when applying the data scaling law in graph classification tasks. To support this argument, we demonstrate that models trained on datasets with identical graph counts but different total numbers of edges exhibit distinctive scaling behaviors. (ii) We further extend the data scaling law to node classification and link prediction tasks, demonstrating the generality of using the number of edges as the data metric.
2 Related Work
Functional forms of neural scaling laws. Neural scaling laws give the quantitative relationship between the model performance and three main factors: the number of model parameters $N$, the size of the training dataset $D$, and the amount of computation $C$. The relationship is generally described as a power-law, which has widely been applied in the domains of image classification (Hestness et al., 2017; Henighan et al., 2020; Sharma & Kaplan, 2022; Zhai et al., 2022), generative language modeling (Kaplan et al., 2020; Henighan et al., 2020; Hoffmann et al., 2022), and natural language translation (Gordon et al., 2021; Ghorbani et al., 2021; Bansal et al., 2022; Fernandes et al., 2023).
Using the test error $\epsilon$ as a metric for model performance, Hestness et al. (2017) proposes the basic functional form of neural scaling laws as follows:
$\epsilon=aX^{-b}+\epsilon_{\infty}$ (1)
The variable $X$ represents any of the parameters $N$, $D$, or $C$. The positive constants $a$ and $b$ govern the rate of error reduction as a function of $X$, while $\epsilon_{\infty}>0$ specifies the minimum achievable error. Equation 1 is appropriate for regression tasks but is ill-defined for classification tasks, since the error rate $\epsilon_{r}$ remains finite for models with zero training. To fix this problem, Zhai et al. (2022) proposes a modified function:
$\epsilon_{r}=a(X+c)^{-b}+\epsilon_{\infty}$ (2)
where a positive shift coefficient $c$ is added to avoid having infinite $\epsilon_{r}$ when $X$ equals to zero. In this work, we will base our following investigations of graph classification tasks on this equation.
The relationship between $N$, $D$ and overfitting. The negative exponent $-b$ in Equation 2 indicates that the performance gain will be diminishing if only one of the parameters $N$ and $D$ is increased. To break this bottleneck, the dataset size and the model size need to be enlarged together. Specifically, Kaplan et al. (2020) find that the model size and dataset size should increase together as $D\sim N^{\alpha}$ ($10<\alpha<$ 1) to attain optimal return for compute budgets. It is also observed that overfitting will likely occur if $N$ surpasses $D^{1/\alpha}$ significantly, which leads to negative performance gain.
Graph learning and scaling. The graph is unique in representing non-Euclidean data with irregular and flexible structures, e.g., social networks (Yanardag & Vishwanathan, 2015), molecules (Gilmer et al., 2017), and protein networks (Szklarczyk et al., 2019). Despite the popularity of graph machine learning, neural scaling laws on graphs remain underexplored. Some works have looked into the neural scaling laws for molecule representation learning (Chen et al., 2023), but the exploration is limited to the data and model types. It is unclear how the conclusion can be generalized to other graphs. To the best of our knowledge, we are the first to investigate neural scaling laws for the general graph domain.
3 Preliminaries
(a) Model scaling on pcqm4mv2, $R^{2}=0.97$.
(b) Data scaling on pcqm4mv2, $R^{2}=0.98$.
Figure 1: The model and data scaling behaviors of GIN on pcqm4mv2. For model scaling, the color corresponds to the training data fraction; for data scaling, the color corresponds to the number of model parameters. The lower the mean square error is, the better the model performance is. The high $R^{2}$ values indicate that the scaling law curves fit well with the empirical values.
(a) Model scaling on ogbg-ppa, $R^{2}=0.98$.
(b) Data scaling on ogbg-ppa, $R^{2}=0.98$.
Figure 2: The model and data scaling behaviors of GIN on ogbg-ppa. For model scaling, the color corresponds to the training data fraction; for data scaling, the color corresponds to the number of model parameters. The higher the accuracy is, the better the model performance is. The high $R^{2}$ values indicate that the scaling law curves fit well with the empirical values.
Orgnizations. The main focus of this work is on the two basic forms of neural scaling laws for the graph domain, i.e., the modeling scaling law and the data scaling law. To consolidate our conclusion, our investigations focus on the well-studied supervised graph learning while leaving the under-studied unsupervised learning tasks (Sun et al., 2022) as one future work. The following sections will be arranged as follows: Section 4 verifies the validity of the basic forms of neural scaling laws on the graphs. Section 5 investigates deeper into the model scaling law of graph models. We observe the model scaling collapse phenomenon when graph models are excessively large and identify overfitting as the potential reason. Moreover, we reveal that model depth can affect the model scaling behaviors, which is a unique phenomenon in graph domains. In Section 6, we propose the number of edges is a suitable data metric when applying the data scaling law on graphs. To support the argument, we exhibit that models would have distinctions in scaling behavior when trained on datasets with different numbers of edges. Furthermore, we present that using the number of edges as the data metric could provide a unified data scaling law for various graph learning tasks, e.g., node classification, link prediction, and graph property prediction.
[…]
$s=-a(X+c)^{b}+s_{\infty}$ (3)
where $s$ is the test score corresponding to specific metrics and $s_{\infty}=1-\epsilon_{\infty}$. Equation 3 will be used in the following sections to fit the scaling law on classification tasks. The details of datasets and experimental settings can be found in Appendix A.
4 The Validity of Basic Scaling Laws on Graphs
(a) Scaling behaviors on pcqm4mv2, $R^{2}=0.98$.
(b) Scaling behaviors on ogbg-ppa, $R^{2}=0.86$.
Figure 3: The comprehensive scaling behaviors of GIN on pcqm4mv2 and ogbg-ppa datasets. For pcqm4mv2, the color corresponds to the value of mean square error; for ogbg-ppa, the color corresponds to accuracy. The surfaces predicted by the neural scaling law fit well with the empirical values on both datasets, as indicated by the high $R^{2}$ values.
We start our analysis by examining the basic forms of neural scaling laws in the graph domain. In other words, we aim to explore how the model performance changes with the dataset size and model size. Hence, we first evaluate the performance of the models across various sizes and training dataset scales. To simplify the experimental settings, we vary the model size by increasing the model width while fixing the model depth, and the effect of model depth will be further explored in Section 5. To obtain subsets with different sizes, we randomly subsample the original training set by different ratios. Then, we fit the scaling law Eq. 1 and 3 to the empirical performance data using the least square method. To measure the fitting quality, the coefficient of determination $R^{2}$ is adopted as a metric. In this subsection, we focus on GIN and leave the results of GCN in Appendix B. Our experiments are conducted on two large graph property prediction datasets: pcqm4mv2 and ogbg-ppa.
We conduct experiments for both model scaling and data scaling behaviors on the graph regression dataset pcqm4mv2. For regression tasks, the performance is measured by mean square error (MSE), hence we use Eq. 1 to fit the empirical MSEs. For model scaling, we fix the number of layers for the model to $6$ and vary the model width from $64$ to $1,024$. The experiments are repeated for different data fractions chosen from $\left[0.1,0.2,0.3,0.4,0.5,0.7,1\right]$. The results, shown in Figure 1(a), depict each model’s MSE as dots, with dashed lines representing fitted scaling law curves and colors indicating specific training data fractions. We make the following observations from Figure 1(a): (1) the MSE tends to decrease when increasing the model size and fixing the dataset size. (2) the MSE gradually decreases and approaches a lower bound $\epsilon$ asymptotically as the model becomes larger. The scaling curves corresponding to larger dataset sizes would have smaller $\epsilon$, i.e., higher upper bound of the model performance. The observations in Figure 1(b) exhibit similar patterns when scaling up data size while fixing the model size. Generally speaking, both model scaling curves in Figure 1(a) and data scaling curves in 1(b) fit well to empirical values with high $R^{2}$s, indicating that Eq. 1 could be leveraged to describe the neural scaling laws for graph regression tasks.
The scaling behaviors on the ogbg-ppa dataset are illustrated in Figure 2. The way of varying dataset sizes and model sizes is similar to that of pcqm4mv2. However, since the task on ogbg-ppa is graph classification, we use accuracy as the performance metric and apply Eq. 3 to fit the scaling behaviors. Observations from Figure 2 align with those from Figure 1, suggesting that Equation 3 can describe the scaling behavior for graph classification tasks.
Moreover, similar to Rosenfeld et al. (2019), the model scaling law and data scaling law formulations can be combined into a comprehensive form for regression tasks:
$\epsilon=a_{1}D^{b_{2}}+a_{2}N^{b_{2}}+\epsilon_{\infty}$ (4)
For classification tasks, the comprehensive form is:
$s=-a_{1}(D+c_{1})^{b_{1}}-a_{2}(N+c_{2})^{b_{2}}+s_{\infty}$ (5)
We verify the comprehensive forms by fitting the empirical values to Equation 4 and 5. The experimental results are shown in Figure 3, where the high $R^{2}$ values suggest that the model scaling and data scaling laws can be well unified by Equation 4 and 5.
Observation 1. The model and data scaling behaviors in the graph domain can be described by neural scaling laws.
The above observations suggest that the scaling behaviors for graph property prediction tasks follow neural scaling laws and can be correctly predicted by Eq. 1 and 3. Thus, it is possible to attain expected performance on graph tasks by enlarging the model and dataset size, exhibiting the potential of building large graph models.
(a) Model scaling on ogbg-molpcba
(b) Training losses of models
(c) Validation losses of models
Figure 4: (a) The empirical performance of larger models falls below the prediction of model scaling law. (b)(c) The training losses and validation losses of models with varying sizes. The color corresponds to the number of model parameters. The model with larger sizes suffers more serious overfitting
Figure 5: More training data can mitigate the overfitting problem. The model size is fixed, while the two curves correspond to different data fractions of the original ogbg-molpcba training set.
5 Graph-specific Phenomena for Model Scaling Laws
In this section, we will dive deeper into specific phenomena of model scaling on graphs. In Section 5.1, we observe the collapse of model scaling in certain situations and discuss the potential reason. In Section 5.2, we identify that models with different depths could exhibit distinct model scaling behaviors. These observations enable us to have a better understanding of the uniqueness of model scaling law on graphs, providing important instructions for developing large graph models.
5.1 Model Scaling Collapse
Despite the overall success, the scaling law could fail in certain situations. In particular, we find that the model scaling law could fail when the model size is too large relative to the training set size, causing the phenomenon called model scaling collapse. To illustrate this phenomenon, we consider an example of scaling up 5-layer GIN on ogbg-molpcba dataset. Similar to the setting in Section 4, we vary the model size by changing the model width while fixing the training set size. Figure 4(a) displays the empirical accuracy of models across sizes alongside the model scaling curve. At small model sizes, the accuracy values fit with the model scaling curve well. However, when the number of parameters is around $10^{7}$, the empirical values start dropping. Thus, the collapse of performance would make the predictions of the scaling law fail, which is referred to as model scaling collapse.
The model scaling collapse poses difficulty in building large graph models following the model scaling law. Thus, it is essential to explore why the collapse happens and how to mitigate it. In this work, we conjecture that one of the potential reasons for the scaling law collapse is overfitting. To dive deeper into the problem, Figure 4(b) and 4(c) present the training loss and validation loss curves, respectively. All the training losses in Figure 4(b) keep dropping in the training process. The larger the model is, the smaller the corresponding training loss will be. Nonetheless, the validation losses of models with more than $10^{7}$ parameters start increasing after several epochs as shown in Figure 4(c), which indicates overfitting. Moreover, the models whose performance falls behind the scaling law predictions suffer more serious overfitting. For instance, models of 7.4M, 13M, and 20M parameters overfit the most in Figure 4(c), and they are the three models on which model scaling law fails in Figure 4(a). Based on these observations, overfitting can be one potential reason for the model scaling collapse. For more results between the model size and overfitting, please refer to Appendix C.
(a) Model scaling of GCN with different layers.
(b) Model scaling of GIN with different layers.
Figure 6: Model scaling behaviors of GNNs with varying model depths. The color corresponds to the model layer. Both GIN and GCN exhibit distinct scaling behaviors with different model depths.
(a) Model scaling of SAT with different layers.
(b) Model scaling of GraphGPS with different layers.
Figure 7: Model scaling behaviors of graph transformers with varying model depths. The color corresponds to the model layer. Both SAT and GraphGPS exhibit distinct scaling behaviors with different model depths.
To eliminate the effect of model scaling collapse, it is crucial to overcome the overfitting. Kaplan et al. (2020) suggest that overfitting would happen if the model size is too large for the training set. To investigate whether more training data can mitigate overfitting, we train a model of fixed number parameters on different fractions of the training set and illustrate the validation loss curves in Figure 5. The model trained on $10$ percent of the original ogbg-molpcba dataset suffers serious overfitting as the validation loss curve suggests. However, the overfitting is significantly mitigated when the data fraction is raised to $50$ percent, which indicates the effectiveness of enlarging the training set. Another evidence is that we have not observed overfitting or model scaling collapse in models of similar sizes ($10^{6}$ parameters) trained on larger datasets, i.e., pcqm4mv2 and ogbg-ppa in Section 4.
Observation 2. The model scaling collapses in the graph domain often happen due to the overfitting.
In graph tasks, the model scaling collapse often occurs with overfitting, which is because the graph dataset sizes are much smaller compared to those in CV or NLP. Overall, the model scaling collapse in graph tasks suggests that the current graph datasets are small for the deep graph models, leaving the models data hungry. Furthermore, it indicates that a crucial step toward the large graph model is to enlarge the size of datasets.
5.2 The Effect of Model Depth
Another unique question in the graph domain is whether the model depth would affect the model scaling law, or more specifically, the scaling law coefficients in Equation 3. Kaplan et al. (2020) find that in the NLP domain, the model depth does not have effects on the scaling coefficients, i.e., the scaling curves for transformers of different numbers of model layers will overlap. Therefore, conducting preliminary experiments to determine the optimal number of layers for large language models (LLMs) is not required, which effectively narrows the hyperparameter search space. Since the deep graph models have different architectures from vanilla transformers, it is desired to study whether the above conclusion still holds.
(a) ogbg-molhiv
(b) reddit-multi-12k
Figure 8: Performance curves of GIN with varying sizes trained on two subsets. The simple subset and the complex subset have the same number of graphs but different numbers of edges. The shade corresponds to the performance variance of repeated experiments.
Figure 9: Performance curves of GIN of varying sizes trained on two subsets. The simple subset and the complex subset have the same number of edges but different numbers of graphs. The performance differences between the two curves are much smaller than the curves in Figure 8(a), suggesting that the number of edges is a better data metric than the number of graphs.
To ensure the generalization of our investigations, we choose two representative GNNs (GCN & GIN) and two powerful graph transformers (SAT & GraphGPS). For these models, we choose the layer number in $[3,4,5]$ and fix it but change the model width to attain varying model sizes. All the models are trained on the whole ogbg-ppa training set and then evaluated for model performance. The model scaling curves of models with different model depths are presented in Figure 6 and 7. Observations can be made as follows: (1) The scaling behaviors of models with different numbers of layers are distinct. For each model, the shape of the scaling law curves would change when the model depth changes, exhibiting different increasing speeds and upper bounds for performance. (2) The optimal number of layers for different models could be different. For GCN and GraphGPS, models with $3$ layers have the best performance, but for GIN the optimal layer number is $4$, and for SAT is $5$. Similar observations on more datasets can be found in Appendix D.
Observation 3. Deep graph models with varying depths will have different model scaling behaviors.
Current observations support that there is no one-size-fits-all model depth for existing graph learning models. The optimal number of model layers could be task-specific and model-specific. We emphasize the necessity of preliminary experiments for the optimal number of model layers before scaling up the model parameters. Notably, we do not have any principle guidance helping to automatically select the best model layer so far. Such mystery could be attributed to the complicated non-parametric aggregation design in different GNNs with various expressiveness (Anonymous, 2024) and inductive bias (Battaglia et al., 2018). Therefore, it is likely that there is no principle across models. Instead, there could be some rules specific to tasks and models without a consensus. Therefore, such analysis is out of scope targeting a general principle across different model architectures and different tasks. We leave the deeper investigation on it as the future work.
6 Reforming Data Scaling Law on Graphs
In this section, we address the irregular-size problems of graphs for the data scaling law. Specifically, we re-examine data metrics in previous studies and reveal the disadvantage of the number of graphs as the data metric in the data scaling law. Instead, we identify the total number of edges in the training set as a better metric with experimental evidence. Furthermore, the data scaling law with the number of edges as the data metric can be naturally extended to node classification and link prediction tasks, potentially providing a unified view of data scaling behaviors in the graph domain.
(a) Node classification on ogbn-arxiv, $R^{2}=0.97$.
(b) Link prediction on ogbl-collab, $R^{2}=0.99$.
Figure 10: Both node classification and link prediction follow the data scaling law with the total edge number as the data metric.
6.1 Number of Edges as Data Metric
In previous sections, the data metric in scaling laws is the subsampling ratio of total graphs, i.e., how many graphs in the whole training set are used (Rosenfeld et al., 2019; Chen et al., 2023). Nonetheless, using the number of graphs as the metric neglects the size irregularity of graphs. For instance, a graph with $10$ edges and a graph with $1,000$ edges are treated equivalently when the number of graphs is the metric, even though they could differ remarkably in training costs and impacts on model performance improvement.
To illustrate the disadvantages of adopting the number of graphs as the data metric, we first divide the training sets of one dataset into two subsets as follows: (1) We first arrange all graph samples of class $i$ into a list $\mathcal{G}_{i}$ in an ascending order of their numbers of edges. (2) We choose the first half of all $\mathcal{G}_{i}$ as the “simple” training subset, and then the second half as the “complex” subset. Thus, these two subsets have the same number of graphs while the graphs in the complex subset have more edges. The validation and test sets remain as the original. Then we train GIN of varying sizes on these two subsets and report the performance on ogbg-molhiv and reddit-multi-12k in Figure 8. Results on more datasets can be found in Appendix E.
If the number of graphs can accurately metric the data volume, the curves of simple and complex subsets in Figure 8 should have small differences or even overlap with each other. However, their performance curves exhibit considerable differences, indicating that the two subsets contain different amounts of data. For comparison, we attain the two subsets in another way: the simple subset has the first $m$ samples in $\mathcal{G}_{i}$, and the complex subset has the remaining $n$ samples, where the two subsets contain the same number of edges. It is easy to validate that $m>n$. Then we train models of varying sizes on these two subsets and report their performance on ogbg-molhiv in Figure 9. Compared to Figure 8(a), the curves in Figure 9 have much smaller differences, which indicates that the total number of edges could be a better metric than the number of graphs.
Observation 4. The number of edges is a better data metric for data scaling law compared to the number of graphs.
In addition, using the total number of edges as the data metric in scaling laws has two more advantages: (1) Knowing the GNN architecture and the number of edges in a dataset, the compute cost of training can be directly calculated, which enables us to build connections between data scaling law and compute scaling law as in Kaplan et al. (2020). The detailed explanation is in Appendix F. (2) Furthermore, using the number of edges as the data metric potentially provides a unified way to describe the data scaling behaviors of node classification and link prediction tasks, which will be thoroughly studied in the next subsection.
6.2 Results on Node Classification and Link Prediction
There is only one graph in most of the node classification and link prediction datasets, thus it is infeasible to use the number of graphs to measure the training data volume. Nonetheless, the total number of edges is able to serve as the data metric for these tasks. Moreover, with the total number of edges as the data metric, we can potentially attain a unified data scaling law on graphs by directly extending Eq. 3 to node classification and link prediction tasks. To verify this, we conduct experiments on the data scaling behaviors on both node classification and link prediction tasks. For node classification, we train a $3$-layer GraphSAGE (Hamilton et al., 2017) on ogbn-arxiv, where the hidden vector size is set to $128$. For link prediction, we train a $3$-layer SEAL (Zhang et al., 2021) on ogbl-collab, where the model takes one-hop neighbors of the target node pairs as input and its hidden vector size is set to $256$. We take the total number of edges on which models are trained as the training set size for both tasks. The results are demonstrated in Figure 10. Similar to graph classification tasks, Eq. 3 fits well to the data scaling behaviors, as evidenced by the high $R^{2}$ values.
Observation 5. Node classification, link prediction, and graph classification tasks could follow the same form of data scaling law with the number of edges as data metric.
7 Conclusions
In this work, we take the first step to investigate neural scaling laws on graphs. In particular, we focus on two basic forms of neural scaling laws: the model scaling law and the data scaling law. We first verify the formulations that can describe both scaling behaviors in the graph domain. For the model scaling law, we identify and investigate some graph-specific phenomena, revealing different observations from CV and NLP domains. For the data scaling law, we reform it by adopting the number of edges as the data metric and extend the reformed version to node classification and link prediction tasks. Our findings shed light on the scaling behaviors of deep graph models and can provide valuable instructions for developing large graph models.
Limitations. Next, we discuss the limitations of this work, which could be further explored as future works. (1) There is no large graph dataset of similar sizes as the large CV or NLP datasets. We conduct our experiments on the largest academic graph datasets that we can attain. (2) Since there has not been a well-established pretraining task for graph learning, we limit our investigation to the supervised learning on graphs. (3) In our exploration, we assume that all node (edge) features are in a unified feature space. We do not consider heterogeneous feature problems in this work.
Impact Statements
In this paper, we investigate neural scaling laws on graphs, which describe how the model performance will grow with increasing model and data sizes. Neural scaling laws can provide instructions for building large graph models, predicting the costs of computing and data labeling to achieve the expected performance. Moreover, it may incentivize more investments in data collecting and even the construction of larger graph datasets for various tasks. We anticipate that our findings will serve as essential guidance for the development of large graph models.
Appendix A Emperiment Settings.
Here we introduce the training datasets used and training hyperparameters in the experiments.
Datasets. The basic situations of the datasets we use are listed in the Tabel 1 and  2 below.
Table 1: Graph Property Prediction Datasets Statistics.
pcqm4mv2 ogbg-ppa ogbg-molhiv ogbg-molpcba reddit-multi-12k
Domains Molecule Protein Molecule Molecule Social
#Graphs 3,746,619 158,100 41,127 437,929 11,929
#Total Nodes 52,970,652 38,481,540 1,048,739 11,386,154 4,669,130
#Total Edges 54,546,813 358,270,410 1,130,993 12,305,805 5,450,241
Tasks Regression Classification Classification Classification Classification
Metrics MSE Accuracy ROC-AUC AP Accuracy
Table 2: Node Classification and Link Prediction Datasets Statistics.
ogbn-arxiv ogbl-collab
Domains Citation Social
#Nodes 169,343 235,868
#Edges 1,166,243 1,285,465
Tasks Node Classification Link Prediction
Metrics Accuracy Hits@20
Training Details. For GIN and GCN models, we set the initial learning rate to 0.001 and use Adam optimizer (Kingma & Ba, 2014) in the training process. For SAT and GraphGPS, we directly adopt the hyperparameters and optimizers that the authors present in their GitHub repos. For ogbg-ppa, ogbn-arxiv, and ogbl-collab datasets, the training epoch number is 50. For pcqm4mv2, ogbg-molhiv, ogbg-molpcba, and reddit-multi-12k datasets, the training epoch number is 100. We choose the test performance with the best corresponding validation performance as the final performance of the model.
Appendix B More Results on the Validity of the Neural Scaling Law on Graphs
Here we show the scaling behaviors of GCN on the pcqm4mv2 and ogbg-ppa datasets. All the empirical values fit well with the neural scaling law.
(a) Model scaling on pcqm4mv2, $R^{2}=0.98$.
(b) Data scaling on pcqm4mv2, $R^{2}=0.98$.
Figure 11: The model and data scaling behaviors of GCN on pcqm4mv2. For model scaling, the color corresponds to the training data fraction; for data scaling, the color corresponds to the number of model parameters. The lower mean square error is, the better model performance is. The high $R^{2}$ values indicate that the scaling law curves fit well with the empirical values.
(a) Model scaling on pcqm4mv2, $R^{2}=0.98$.
(b) Data scaling on pcqm4mv2, $R^{2}=0.99$.
Figure 12: The model and data scaling behaviors of GCN on ogbg-ppa. For model scaling, the color corresponds to the training data fraction; for data scaling, the color corresponds to the number of model parameters. The higher performance is, the better model performance is. The high $R^{2}$ values indicate that the scaling law curves fit well with the empirical values.
(a) Scaling behaviors on pcqm4mv2, $R^{2}=0.97$.
(b) Scaling behaviors on ogbg-ppa, $R^{2}=0.97$.
Figure 13: The comprehensive scaling behaviors of GCN on pcqm4mv2 and ogbg-ppa datasets. For pcqm4mv2, the color corresponds to the value of mean square error; for ogbg-ppa, the color corresponds to accuracy. The surfaces predicted by the neural scaling law fit well with the empirical values on both datasets, as indicated by the high $R^{2}$ values.
Appendix C More Results on the Model Scaling Collapse
Here we present more results on the relationship between the model size and the degree of overfitting. In general, the observation in the below figures indicate that the lager model is, the more serious it would overfit.
Figure 14: The training losses and validation losses of GCN with varying sizes on Code2 Dataset. The color corresponds to the number of model parameters. The figure shows that the model with larges size has more serious overfitting.
Figure 15: The training losses and validation losses of GIN with varying sizes on Code2 Dataset. The color corresponds to the number of model parameters. The figure shows that the model with larges size has more serious overfitting.
Figure 16: The training losses and validation losses of GCN with varying sizes on ogbg-ppa Dataset. The color corresponds to the number of model parameters. The figure shows that the model with larges size has more serious overfitting.
Appendix D More Results on the Impacts of Model Depths
We present the experiment results of model scaling behaviors with varying model depths on more datasets. In the following figures, models with different depths exhibit distinct model scaling behaviors.
(a) Model scaling of GCN with different layers.
(b) Model scaling of GIN with different layers.
Figure 17: Model scaling behaviors of GNNs with varying model depths on reddit-multi-12k. The color corresponds to the model layer. Both GIN and GCN exhibit distinct scaling behaviors with different model depths.
(a) Model scaling of GCN with different layers.
(b) Model scaling of GIN with different layers.
Figure 18: Model scaling behaviors of GNNs with varying model depths on ogbg-molhiv. The color corresponds to the model layer. Both GIN and GCN exhibit distinct scaling behaviors with different model depths.
(a) Model scaling of GCN with different layers.
(b) Model scaling of GIN with different layers.
Figure 19: Model scaling behaviors of GNNs with varying model depths on ogbg-molpcba. The color corresponds to the model layer. Both GIN and GCN exhibit distinct scaling behaviors with different model depths.
Appendix E More Results of the Disadvantage of the number of Graphs as Data Scaling Metric
Here we present more results of the experiments in Section 6.1. The settings are the same as in Section 6.1, except we replace GIN with GCN. From the below figure, we observe that performance curves have obvious differences when two subsets have the same number of graphs. But when the two subsets have the same number of edges, the differences between the two curves become much smaller, supporting that the number of edges can better measure the data volume.
Figure 20: Performance curves of GCN of varying sizes trained on two subsets. In (a), the simple subset and the complex subset have the same graph number but different edge numbers. In (b), the simple subset and the complex subset have the same edge number but different graph numbers. The shade corresponds to the performance variance of repeated experiments. This figure supports the same conclusion as in Section 6.1.
Appendix F The Relationship between Edge Number and FLOPs
Here we show that the training compute can be calculated when we know the total edge number and the model architecture. We consider the general Message Passing Neural Networks, which can be formulated as above:
$m^{t+1}_{v}=U(h^{t}_{v},\sum_{w\in N(v)}M(h^{t}_{v},h^{t}_{w},e_{vw}))$ (6)
where $U$ and $M$ can be assumed to contain $x$ and $y$ add-multiply operations, respectively. Suppose the total node number in the dataset is N, the total edge number is E, and the model has L layers. For node $i$, aggregating messages from its neighbors needs $x+yd_{i}$ add-multiply operations, where $d_{i}$ is the degree of node $i$. For the whole datasets, there would be $\sum_{i}x+yd_{i}\approx 2yE$. For the whole model, one feed-forward pass needs $2yEL$ add-multiply operations. Thus, we have shown that we can calculate the traing FLOPs if we know the total edge number in the dataset.
Appendix G The Model Scaling Behaviors of Node Classification and Link Prediction
Here we investigate the model scaling behaviors of node classification and link prediction tasks. The observations in the below figure show that SEAL trained on ogbl-collab has obvious performance gains when the model size is increased, while this is not the case for GraphSAGE on ogbn-arxiv. We leave further exploration of these topics for future works.
(a) Model scaling behaviors on ogbn-arxiv.
(b) Model scaling on ogbl-collab.
Figure 21: In (b), the model performance has obvious gains on the link prediction task when the hidden vector size is increased, while it is not for the node classification task in (a).